Please help me to solve the following problems:
Solve each proportion:
34) b/3=-3/4

Use a proportion to solve each problem:
44) Fast Food. If four out of five doctors prefer fast food, thn ata convention of 445 doctors, how many prefer fast food?
Thank you very much-

Answer by checkley75(3666)

B/3=-3/4 NOW WE NEED TO CROSS MULTIPLY. THUS:
4B=-3*3
4B=-9 NOW WE DIVIDE BOTH SIDES BY 4. THUS:
B=-9/4
B=-2.25
PROOF
-2.25/3=-3/4
-2.25*4=-3*3
-9=-9